# Dissecting Lepton Number Violation in the Left-Right Symmetric Model: $0\nu\beta\beta$ decay, Møller scattering, and collider searches

### v1

@@ III Constraints and prospects at high-energy colliders
For LEP, we consider an average center-of-mass energy of 195.6 GeV with an integrated luminosity of 745 pb-1 Schael et al. (2013). We will focus on the Circular Electron Positron Collider (CEPC) Cheng et al. (2022) at 240 GeV with an integrated luminosity of 20 ab-1, and the Future Circular Collider (FCC-ee) Agapov et al. (2022) at 240 GeV with an integrated luminosity of 5.1 ab-1.
IV Constraints and prospects in low-energy experiments
In this section, we will investigate the sensitivities to the LRSM in the low-energy experiments. The presence of right-handed charged currents impacts the mixings of $K$ and $B$ mesons, which have been investigated in Refs. Bertolini et al. (2014); Dekens et al. (2021); Zhang et al. (2008). However, the resulting constraints on the mass of $W_{R}$ are weaker than that from direct searches at the LHC Li et al. (2021); Aad et al. (2023a). The experimental constraint on $M_{W_{R}}$ in Sec. III can be directly translated into the equivalent bound $\lambda<1.58\times 10^{-4}$ using Eq. (20). Additionally, $\xi<0.8$ is required to ensure the validity of perturbative expansions in the LRSM scalar sector Maiezza et al. (2010); Nemevšek and Zhang (2024).
The left-right mixing angle $\zeta$ is related to the parameters $\xi$ and $\lambda$ as shown in Eq. (21). It is important to highlight that to resolve the deviation from the CKM unitarity using the determination of $|V_{ud}|$ from the nuclear superallowed $0^{+}\to 0^{+}$ decay Seng et al. (2018, 2019); Czarnecki et al. (2019), the left-right mixing angle $\zeta\in[4.5,11]\times 10^{-4}$ is requisite Dekens et al. (2021). This leads to an apparent incompatibility between the CKM and collider implications. However, for the scope of this analysis, we will refrain from delving into the resolution of the discrepancy, and we will adopt the collider bounds.
Besides, $\xi$ and $\lambda$ can be also simultaneously constrained by the $\rho$ parameter Czakon et al. (1999)
$\rho=1+\left[-\frac{2\xi}{1+\xi^{2}}+\frac{(1-\tan^{2}\theta_{W})^{2}}{2}\right]\lambda\,,$ (23)
@@ IV Constraints and prospects in low-energy experiments
In our analysis, we choose different benchmark values consistent with the aforementioned constraints: (I) $M_{W_{R}}=7\,\mathrm{TeV}$, $\xi=0.35$, (II) $M_{W_{R}}=15\,\mathrm{TeV}$, $\xi=0.75$, and (III) $M_{W_{R}}=25\,\mathrm{TeV}$, $\xi=0.8$. The benchmarks for $\xi=0$ and $M_{W_{R}}=7$, $15$ and $25\,\mathrm{TeV}$ are also considered.
@@
As mentioned in Sec. I, searches for CLFV in $\mu\to e\gamma$, $\mu\to eee$, and $\mu\to e$ conversion in nuclei have set stringent constraints on physics beyond the standard model Lindner et al. (2018). The constraints on the left-handed doubly-charged scalar $\delta_{L}^{\pm+}$ are Dev et al. (2018); Akeroyd et al. (2009)
$\mu\to eee:\ M_{\delta_{L}^{\pm+}}\left|\left(f_{L}\right)_{ee}^{\dagger}\left(f_{L}\right)_{e\mu}\right|^{-1/2}>208\,\mathrm{TeV}\;,$ (24)
$\mu\to e\gamma:\ M_{\delta_{L}^{\pm+}}\left|\sum_{\ell}\left(f_{L}\right)_{\mu\ell}^{\dagger}\left(f_{L}\right)_{e\ell}\right|^{-1/2}>65\,\mathrm{TeV}\;,$
@@
where $M_{\delta_{L}^{\pm+}}$ denotes the mass of $\delta_{L}^{\pm+}$. The right-handed doubly-charged scalar $\delta_{R}^{\pm+}$, however, could evade these constraints if the left-right symmetry is not assumed in the Yukawa sector, namely $\mathbf{f_{R}}\neq\mathbf{f_{L}}$, which can be realized by extending the minimal LRSM Chang et al. (1984); Fukuyama et al. (2022).
IV.1 Neutrinoless double beta decay
In this subsection, we calculate the half-life of ${0\nu\beta\beta}$ decay using the EFT framework. In this approach, the heavy degrees of freedom are integrated out at the right-handed scale ($M_{W_{R}}$) and matched to all effective operators that are invariant under the SM gauge group containing only SM fields, the so-called standard model EFT (SMEFT). The higher-dimensional operators then evolve to the electroweak scale ($M_{W}$), where the remaining heavy fields are integrated out again. At even lower energies, around $\Lambda_{\chi}\sim\rm GeV$, the resulting EFT operators are mapped onto hadronic LNV operators by the use of chiral EFT, the low-energy EFT of QCD.
The relation in Eq. (12) indicates that the mass of the right-handed neutrino is correlated with the magnitude of the Yukawa coupling, $\widehat{M}_{N}\sim{\bf f_{R}}\,v_{R}$, where the indices are omitted for simplicity. Exploring small couplings would imply light right-handed neutrinos. If $m_{4,5,6}\buildrel<\over{{}_{\sim}}\Lambda_{\chi}$, the right-handed neutrino fields should be kept as dynamical degrees of freedom in the approach of neutrino-extended SMEFT ($\nu$SMEFT). We adopt the framework developed in Refs. Dekens et al. (2020); de Vries et al. (2022) to study the contributions to the half-life of ${0\nu\beta\beta}$ decay from right-handed neutrinos with arbitrary mass.
Figure 1: Feynman diagrams contributing to ${0\nu\beta\beta}$ decay in the LRSM. The cross vertices $(\bm{\times})$ in between $W_{L}-W_{R}$ propagators denote the $W_{L}-W_{R}$ mixing. In the limit of type-II dominance, there is no mixing between left-handed $\nu$ and right-handed $N$ neutrinos.
The Feynman diagrams for the contributions to ${0\nu\beta\beta}$ decay in type-II dominance scenario are shown in panels (a)-(d) of Fig. 1. We also consider the contributions from right-handed doubly-charged scalar $\delta_{R}^{--}$ in panels (e) and (f). The inverse half-life can be expressed as
$\left(T^{0\nu}_{1/2}\right)^{-1}=$ $g_{A}^{4}\Big{[}G_{01}\left(|\mathcal{A}_{L}|^{2}+|\mathcal{A}_{R}|^{2}\right)$
@@ IV.1 Neutrinoless double beta decay
where the coupling constant $g_{A}=1.271$, and $G_{01}$, $G_{04}$ denote the phase-space factors Cirigliano et al. (2018). The amplitudes $\mathcal{A}_{L,R}$ are defined as
@@
where $m_{i}$ are the active ($i=1,2,3$) and sterile ($i=4,5,6$) neutrino masses in Eq. (6). The subamplitudes $\mathcal{A}_{L,R}(m_{i})$ are given by de Vries et al. (2022); Dekens et al. (2020)
@@
Here, $m_{e}$ is the electron mass, and $\mathcal{M}_{A,V}$ are nuclear matrix elements (NMEs). More details can be found in Refs. Prezeau et al. (2003); Cirigliano et al. (2017, 2018); Dekens et al. (2020); de Vries et al. (2022).
@@
in the EFT approach with dimension-9 effective operators Prezeau et al. (2003) are explicitly given in Ref. Cirigliano et al. (2018). Here, $m_{N}\sim 1\,\mathrm{GeV}$ is the nucleon mass. For conciseness, we will only repeat the Wilson coefficients of the relevant dimension-9 effective operators de Vries et al. (2022) as follows:
$C_{1R}^{(9)\prime}(M_{\delta_{R}^{\pm\pm}})$ $=-\dfrac{g^{4}v^{5}}{4M_{W_{R}}^{4}}(V_{ud}^{R})^{2}\dfrac{m_{i}}{M_{\delta_{R}^{\pm\pm}}^{2}}(P_{s}U^{*})^{2}_{ei}\;,$ (37)
@@
The Wilson coefficients $C_{1R}^{(9)\prime}(M_{\delta_{R}^{\pm\pm}})$ and $C_{1R,4R}^{(9)}(M_{W})$ are defined at different scales. The QCD renormalization-group-evolution equations of them and the other Wilson coefficients above are given in Refs. Cirigliano et al. (2018); de Vries et al. (2022). Note that the contributions of right-handed neutrinos with masses $m_{4,5,6}>\Lambda_{\chi}$ can also be integrated out, rendering dimension-9 effective operators Li et al. (2021). The resulting contributions to ${0\nu\beta\beta}$ decay have also been included in Eq. (28) in the framework developed in Refs. de Vries et al. (2022); Dekens et al. (2020).
IV.2 Parity-violating Møller scattering
In Møller scattering Derman and Marciano (1979), parity-violating asymmetries in polarized electron scattering can constrain the 4-electron effective vertex in a model-independent way. For that purpose, the MOLLER Collaboration proposes to measure the asymmetry in the scattering of polarized electrons off unpolarized electrons using the $11\,\mathrm{GeV}$ beam in Hall A at Jefferson Lab to measure the electron’s weak charge $Q_{W}^{e}$ to an overall accuracy of 2.4% Benesch et al. (2014). This measurement could represent an enhancement in fractional precision exceeding five times that of the only other similar one done by the E158 experiment at SLAC Anthony et al. (2005). The MOLLER sensitivity can be expressed as the following lower bound:
@@ IV.2 Parity-violating Møller scattering
[…]
where $\Lambda$ is the mass scale of new physics Benesch et al. (2014); Eichten et al. (1983).
In the context of the LRSM, one can relate parity-violating Møller scattering, which conserves the lepton number, to ${0\nu\beta\beta}$ decay that violates the lepton number by two units. The leptonic Yukawa interactions in Eq. (II) allow for both the left- and the right-handed doubly-charged scalars to contribute to Møller scattering. These contributions consist of an $s$-channel exchange depicted in Fig. 2. The corresponding parity-violating effective Lagrangian can be written as Dev et al. (2018)
[…]
where we have used the Firez identity Nieves and Pal (2004); Liao and Liu (2012)
@@
It is worth noting that there exists an additional contribution to Møller scattering from the $t$-channel exchange of the neutral scalar from the scalar bidoublets. However, the corresponding amplitude does not interfere with the SM amplitude and thus contributes at the $1/\Lambda^{4}$ order, which can be easily verified with the following Fierz identity
@@
In principle, the effective Lagrangian in Eq. (42) has left- and right-handed contributions. However, the severe constraints in Eq. (LABEL:eq:cLFV-constraints) from the searches for $\mu\to eee$ and $\mu\to e\gamma$ imply that both the couplings $(f_{L})_{ee}$ and $(f_{L})_{e\mu}$ are tiny or $\delta^{\pm\pm}_{L}$ is heavy  [^1] The couplings $(f_{L})_{ee}$ and $(f_{L})_{e\mu}$ are correlated with the neutrino masses and PMNS matrix in the type-II seesaw dominance scenario, so that $(f_{L})_{ee}\sim(f_{L})_{e\mu}$ is expected.. Therefore, only the contribution from $\delta_{R}^{\pm\pm}$ will be considered in light of the MOLLER prospect.
@@
We can directly match the corresponding expressions in Eq. (42) with Eq. (40), where the effective mass scale is $\Lambda=M_{\delta_{R}^{\pm\pm}}$ and the effective couplings are $|g_{RR}|^{2}=2|(f_{R})_{ee}|^{2}$ and $g_{LL}=0$. Therefore, the parity-violating precision measurement would constrain the ratio between the right-handed doubly-charged scalar mass $M_{\delta_{R}^{\pm\pm}}$ and its Yukawa coupling to electrons $(f_{R})_{ee}$ Benesch et al. (2014),
@@
at $95\%$ confidence level (CL)  [^2] Notice that Ref. Dev et al. (2018) misses a factor of two in the mapping into the 4-fermion effective interaction in Eq. (41) that is corrected here..
V Results and Discussion
This section discusses in detail the results of ${0\nu\beta\beta}$ decay in the EFT approach and its interplay with the other searches and measurements. In the type-II seesaw scenario of the LRSM, the $6\times 6$ block-diagonal matrix in Eq. (10) is independently diagonalized by the $3\times 3$ mass matrices $M_{\nu}$ and $M_{N}$ corresponding to left-handed (active) and right-handed (sterile) neutrinos, respectively. Notice the right-handed neutrinos could have masses of the order of the right-handed scale $v_{R}$ but can also be significantly lower.
To establish a concrete $f_{R}$-correlation between the observables under consideration, we will assume the neutrino mixing matrix $U_{R}=1$ for simplicity and focus our study on the effects of the right-handed neutrino $\nu_{4}$ with the mass
$M_{N_{R}}\equiv m_{4}=\sqrt{2}\,v_{R}\,(f_{R})_{ee}\,,$ (46)
while the other right-handed neutrinos decouple. On the other hand, from Eq. (11), the masses of the active neutrinos are completely determined once the lightest active neutrino mass is chosen and the hierarchy is specified. The oscillation parameters of active neutrinos encoded in the PMNS matrix are quoted from Ref. de Salas et al. (2021).
We will start by investigating the sensitivity of ${0\nu\beta\beta}$-decay searches. We adopt the central values of the oscillation parameters in Table 3 from Ref. de Salas et al. (2021) and marginalize over the Majorana phases $\lambda_{1,2}$ in calculating the ${0\nu\beta\beta}$-decay rate. The lightest active neutrino mass is randomly generated with the upper bound $m_{\nu_{\min}}<0.05\,\mathrm{eV}$. Hence, the inverse half-life in Eq. (25) depends only on a few parameters
  • The mass of the doubly-charged scalar $M_{\delta_{R}^{\pm\pm}}$ and its Yukawa coupling to electrons $(f_{R})_{ee}$.

### v2

@@ 3 Constraints and prospects at high-energy colliders
For LEP, we consider an average center-of-mass energy of 195.6 GeV with an integrated luminosity of 745 pb-1 [ALEPH:2013dgf]. We will focus on the Circular Electron Positron Collider (CEPC) [CEPCPhysicsStudyGroup:2022uwl] at 240 GeV with an integrated luminosity of 20 ab-1, and the Future Circular Collider (FCC-ee) [Agapov:2022bhm] at 240 GeV with an integrated luminosity of 5.1 ab-1.
4 Constraints and prospects in low-energy experiments
In this section, we will investigate the sensitivities to the LRSM in the low-energy experiments. The presence of right-handed charged currents impacts the mixings of $K$ and $B$ mesons, which have been investigated in Refs. [Bertolini:2014sua, Dekens:2021bro, Zhang:2007da]. However, the resulting constraints on the mass of $W_{R}$ are weaker than those from direct searches at the LHC [Li:2020flq, ATLAS:2023cjo]. The experimental constraint on $M_{W_{R}}$ in Sec. 3 can be directly translated into the equivalent bound $\lambda<1.58\times 10^{-4}$ using Eq. (18). Additionally, $\xi<0.8$ is required to ensure the validity of perturbative expansions in the LRSM scalar sector [Maiezza:2010ic, Nemevsek:2023yjl].
The left-right mixing angle $\zeta$ is related to the parameters $\xi$ and $\lambda$ as shown in Eq. (19). It is important to highlight that to resolve the deviation from the CKM unitarity using the determination of $|V_{ud}|$ from the nuclear superallowed $0^{+}\to 0^{+}$ decay [Seng:2018yzq, Seng:2018qru, Czarnecki:2019mwq], the left-right mixing angle $\zeta\in[4.5,11]\times 10^{-4}$ is requisite [Dekens:2021bro]. This leads to an apparent incompatibility between the CKM and collider implications. However, for the scope of this analysis, we will refrain from delving into the resolution of the discrepancy, and we will adopt the collider bounds.
Besides, $\xi$ and $\lambda$ can also be simultaneously constrained by the $\rho$ parameter [Czakon:1999ga]
$\rho=1+\left[-\left(\frac{2\xi}{1+\xi^{2}}\right)^{2}+\frac{(1-\tan^{2}\theta_{W})^{2}}{2}\right]\lambda\,$, (21)
@@ 4 Constraints and prospects in low-energy experiments
In our analysis, we chose different benchmark values consistent with the aforementioned constraints: (I) $M_{W_{R}}=7\,\mathrm{TeV}$, $\xi=0.35$, (II) $M_{W_{R}}=15\,\mathrm{TeV}$, $\xi=0.75$, and (III) $M_{W_{R}}=25\,\mathrm{TeV}$, $\xi=0.8$. The benchmarks for $\xi=0$ and $M_{W_{R}}=7$, $15$ and $25\,\mathrm{TeV}$ are also considered.
@@
As mentioned in Sec. 1, searches for CLFV in $\mu\to e\gamma$, $\mu\to eee$, and $\mu\to e$ conversion in nuclei have set stringent constraints on physics beyond the standard model [Lindner:2016bgg]. The constraints on the left-handed doubly-charged scalar $\delta_{L}^{\pm\pm}$ are [Dev:2018sel, Akeroyd:2009nu]
[…]
where $M_{\delta_{L}^{\pm\pm}}$ denotes the mass of $\delta_{L}^{\pm\pm}$.
@@
The right-handed doubly-charged scalar $\delta_{R}^{\pm+}$, however, could evade these constraints if the left-right symmetry is not assumed in the Yukawa sector. In the LRSM with $D$-parity breaking, this is evident since parity and the $SU(2)_{\rm R}$ breaking scale decouple from each other [Chang:1983fu], which has been considered in Ref. [Dev:2018sel]. A realistic model [Akhmedov:2024rvp] was recently developed along this line, making it feasible to explore the scenario where $\mathbf{f_{R}}\neq\mathbf{f_{L}}$.
4.1 Neutrinoless double beta decay
Although the literature contains numerous studies regarding $0\nu\beta\beta$} decay in the LRSM (see, for example, Refs [Tello:2010am, Awasthi:2015ota, BhupalDev:2014qbx, Bambhaniya:2015ipg]), the present study includes several novel elements. In particular, results in many of the previous analyses rely upon an expansion of the neutrino propagator $1/(\langle p\rangle^{2}+m_{N}^{2})$ in powers of $\langle p\rangle/m_{N}$, where $\langle p\rangle\simeq 100\,{\rm MeV}$ is the typical momentum transfer and $m_{N}$ is the right-handed neutrino mass. It has been demonstrated that the $0\nu\beta\beta$} decay rate calculated using this approach can benefit from the systematic prescription based on chiral EFT, leading to sizable differences, particularly in cases involving chiral enhancement arising from left-right mixing at the quark level; see, for instance, Refs. [Cirigliano:2018yza, Li:2020flq, Dekens:2020ttz, Cirigliano:2024ccq, Cirigliano:2025vye]. Additionally, the propagator-expansion approach breaks down for light sterile neutrinos, namely $m_{N}\lesssim\langle p\rangle$. Our work is based on a chiral EFT approach [Dekens:2020ttz], which allows us to investigate the effects of a wider range of sterile neutrino masses.
[…]
The Feynman diagrams for the contributions to $0\nu\beta\beta$} decay are shown in panels (a)-(d) of Fig. 1. We also include the contributions from right-handed doubly-charged scalar $\delta_{R}^{--}$ in panels (e)-(g), while those from $\delta_{L}^{--}$ are omitted. Moreover, contributions arising from mixing between left- and right-handed neutrinos are significantly smaller than those from the right-handed neutrino [deVries:2022nyh].
[…]
The inverse half-life can be expressed as
$\left(T^{0\nu}_{1/2}\right)^{-1}=\$$g_{A}^{4}\Big[G_{01}\left(|\mathcal{A}_{L}|^{2}+|\mathcal{A}_{R}|^{2}\right)$
@@ 4.1 Neutrinoless double beta decay
where the coupling constant $g_{A}=1.271$, and $G_{01}$, $G_{04}$ denote the phase-space factors [Cirigliano:2018yza]. The amplitudes $\mathcal{A}_{L,R}$ are defined as
@@
where $m_{i}$ are the active ($i=1,2,3$) and sterile ($i=4,5,6$) neutrino masses in Eq. (6). The subamplitudes $\mathcal{A}_{L,R}(m_{i})$ are given by [deVries:2022nyh, Dekens:2020ttz]
@@
Here, $m_{e}$ is the electron mass, and $\mathcal{M}_{A,V}$ are nuclear matrix elements (NMEs). More details can be found in Refs. [Prezeau:2003xn, Cirigliano:2017djv, Cirigliano:2018yza, Dekens:2020ttz, deVries:2022nyh].
@@
in the EFT approach with dimension-9 effective operators [Prezeau:2003xn] are explicitly given in Ref. [Cirigliano:2018yza]. Here, $m_{N}\sim 1\,\mathrm{GeV}$ is the nucleon mass. For conciseness, we will only repeat the Wilson coefficients of the relevant dimension-9 effective operators [deVries:2022nyh] as follows:
$C_{1R}^{(9)\prime}(M_{\delta_{R}^{\pm\pm}})$ $=-\dfrac{g_{R}^{4}v^{5}}{4M_{W_{R}}^{4}}(V_{ud}^{R})^{2}\dfrac{m_{i}}{M_{\delta_{R}^{\pm\pm}}^{2}}(P_{s}U^{*})^{2}_{ei}\;,$ (35)
@@
The Wilson coefficients $C_{1R}^{(9)\prime}(M_{\delta_{R}^{\pm\pm}})$ and $C_{1R,4R}^{(9)}(M_{W})$ are defined at different scales. The QCD renormalization-group-evolution equations of them and the other Wilson coefficients above are given in Refs. [Cirigliano:2018yza, deVries:2022nyh]. Note that the contributions of right-handed neutrinos with masses $m_{4,5,6}>\Lambda_{\chi}$ can also be integrated out, rendering dimension-9 effective operators [Li:2020flq]. The resulting contributions to $0\nu\beta\beta$} decay have also been included in Eq. (26) in the framework developed in Refs. [deVries:2022nyh, Dekens:2020ttz].
Notice that the CP phase $\alpha$ can contribute to the Wilson coefficients $C_{L}^{(6)}(M_{W_{R}})$ and $C_{1R,4R}^{(6)}(M_{W})$ as a factor of $e^{-{\rm i}\alpha}$ [deVries:2022nyh]. However, the amplitude for $0\nu\beta\beta$ decay depends on the product of these coefficients with the elements of the PMNS matrix. By marginalizing over the Majorana phases, a non-zero value of $\alpha$ will not affect the results of the $0\nu\beta\beta$ decay rate that we have obtained.
4.2 Parity-violating Møller scattering
In Møller scattering [Derman:1979zc], parity-violating asymmetries in polarized electron scattering can constrain the 4-electron effective vertex in a model-independent way. For that purpose, the MOLLER Collaboration proposes to measure the asymmetry in the scattering of polarized electrons off unpolarized electrons using the $11\,\mathrm{GeV}$ beam in Hall A at Jefferson Lab to measure the electron’s weak charge $Q_{W}^{e}$ to an overall accuracy of 2.4% [MOLLER:2014iki]. This measurement could represent an enhancement in fractional precision exceeding five times that of the only other similar one done by the E158 experiment at SLAC [SLACE158:2005uay]. The MOLLER sensitivity can be expressed as the following lower bound:
@@ 4.2 Parity-violating Møller scattering
$\mathcal{L}_{\mathit{eff}}=\sum_{i,j=L,R}\frac{g_{ij}^{2}}{2\Lambda^{2}}(\bar{e}_{i}\gamma^{\mu}e_{i})(\bar{e}_{j}\gamma_{\mu}e_{j})\,$, (39)
where $\Lambda$ is the mass scale of new physics [MOLLER:2014iki, Eichten:1983hw].
In the context of the LRSM, one can relate parity-violating Møller scattering, which conserves the lepton number, to $0\nu\beta\beta$} decay that violates the lepton number by two units. The leptonic Yukawa interactions in Eq. (2) allow for both the left- and the right-handed doubly-charged scalars to contribute to Møller scattering. These contributions consist of an $s$-channel exchange depicted in Fig. 2. The corresponding parity-violating effective Lagrangian can be written as [Dev:2018sel]
${\cal L}_{\rm PV}^{\mathit{eff}}=\frac{|(f_{L})_{ee}|^{2}}{M_{\delta_{L}^{\pm\pm}}^{2}}(\bar{e}_{L}\gamma^{\mu}e_{L})(\bar{e}_{L}\gamma_{\mu}e_{L})~+~(L\leftrightarrow R)\,$, (40)
where we have used the Firez identity [Nieves:2003in, Liao:2012uj]
@@
It is worth noting that there exists an additional contribution to Møller scattering from the $t$-channel exchange of the neutral scalar from the scalar bidoublet. However, the corresponding amplitude does not interfere with the SM amplitude and thus contributes at the $1/\Lambda^{4}$ order, which can be easily verified with the following Fierz identity
@@
Given the severe constraints on the mass of the neutral scalar from the scalar bidoublet [Bertolini:2019out, Dekens:2021bro], their contributions to the Møller scattering can be safely neglected.
In principle, the effective Lagrangian in Eq. (40) has left- and right-handed contributions. However, the severe constraints in Eq. (LABEL:eq:cLFV-constraints) from the searches for $\mu\to eee$ and $\mu\to e\gamma$ imply that both the couplings $(f_{L})_{ee}$ and $(f_{L})_{e\mu}$ are tiny or $\delta^{\pm\pm}_{L}$ is heavy. Therefore, only the contribution from $\delta_{R}^{\pm\pm}$ will be considered in light of the MOLLER prospect.
@@
We can directly match the corresponding expressions in Eq. (40) with Eq. (38), where the effective mass scale is $\Lambda=M_{\delta_{R}^{\pm\pm}}$ and the effective couplings are $|g_{RR}|^{2}=2|(f_{R})_{ee}|^{2}$ and $g_{LL}=0$. Therefore, the parity-violating precision measurement would constrain the ratio between the right-handed doubly-charged scalar mass $M_{\delta_{R}^{\pm\pm}}$ and its Yukawa coupling to electrons $(f_{R})_{ee}$ [MOLLER:2014iki],
@@
at $95\%$ confidence level (CL)  [^3] Notice that Ref. [Dev:2018sel] misses a factor of two in the mapping into the 4-fermion effective interaction in Eq. (39) that is corrected here..
5 Results and Discussion
This section discusses in detail the results of $0\nu\beta\beta$} decay in the EFT approach and its interplay with the other searches and measurements. Notice that the right-handed neutrinos could have masses of the order of the right-handed scale $v_{R}$, but can also be significantly lower. To establish a concrete $f_{R}$-correlation between the observables under consideration, we will focus our study on the effects of the right-handed neutrino $\nu_{4}$ with the mass
$M_{N_{R}}\equiv m_{4}=\sqrt{2}\,v_{R}\,|(f_{R})_{ee}|\,$, (44)
while the other right-handed neutrinos decouple. On the other hand, from Eq. (11), the masses of the active neutrinos are completely determined once the lightest active neutrino mass is chosen and the hierarchy is specified. The oscillation parameters of active neutrinos encoded in the PMNS matrix are quoted from Ref. [deSalas:2020pgw].
We will start by investigating the sensitivity of $0\nu\beta\beta$}-decay searches. We adopt the central values of the oscillation parameters in Table 3 from Ref. [deSalas:2020pgw] and marginalize over the Majorana phases $\lambda_{1,2}$ in calculating the $0\nu\beta\beta$}-decay rate. The lightest active neutrino mass is randomly generated with the upper bound $m_{\nu_{\min}}<0.05\,\mathrm{eV}$. Hence, the inverse half-life in Eq. (23) depends only on a few parameters
  • The mass of the doubly-charged scalar $M_{\delta_{R}^{\pm\pm}}$ and its Yukawa coupling to electrons $|(f_{R})_{ee}|$.